OpenStudy (cherie_magee):

Based on the lesson, which of the following would be the best approach for solving this system by substitution? 5x = y + 6 2x - 3y = 4 Solve the first equation for x. Solve the first equation for y. Solve the second equation for x. Solve the second equation for y.

OpenStudy (radar):

They are all legitimate approaches and would result in solving the problem. However, the second approach would be the simplest.

OpenStudy (radar):

It is simplest, simply because the y is easy to isolate. y=5x-6

OpenStudy (radar):

That can be substituted in the second equation like so: 2x -3(5x-6)=4 2x-15x + 18 = 4 -13x=-18+4 -13x = -14 x=14/13